# Dense Video Object Captioning from Disjoint Supervision

### v2

@@ 3.3 Trajectory captioning
Soft aggregation. A straightforward way to leverage object features over time is to compute a weighted sum to combine them into a single, global trajectory feature. We observe that the association matrix, $\mathbf{A}$ (Sec. References), already serves as a summation weight. Specifically, we set $\mathbf{G}=\frac{\mathbf{A}}{||\mathbf{A}||}\cdot\mathbf{F}$, where $\cdot$ denotes matrix multiplication, and |||||$\cdot$||| | ⋅ | | normalizes $\mathbf{A}$ by rows. Each row of $\mathbf{G}$, $\mathbf{g}_{k}\in\mathbb{R}^{r^{2}\times d}$, therfore denotes an aggregated feature over its trajectory for object $k$.
Hard aggregation. An alternative to weighted temporal averaging is to concatenate and construct new trajectory features. Let $\mathbf{f}_{\tau}=\{\mathbf{f}_{k^{\prime}}\}_{\delta_{k^{\prime}}=\tau}$ be the set of all object features with identity $\tau$. We note $\mathbf{f}_{\tau}$ can be as long as the entire video, and thus it may be expensive to directly use $\mathbf{f}_{\tau}$. Therefore, we uniformly sample a subset of object features from the trajectory, denoted as $\mathbf{g}_{\tau}=\text{UniformSample}(\mathbf{f}_{\delta},m)$, where $\mathbf{g}_{\tau}\in\mathbb{R}^{mr^{2}\times d}$, inspired by [61]. $m$ is the number of sampled frames, and we set $m=6$ following ablations in the supplement.
@@ 3.4 Pretraining with disjoint subtasks
As shown in Fig. 3, our model is trained with the loss function, $L=L_{object}+L_{assoc}+L_{caption}$. When we have full Dense VOC annotations, which supervise each component of our model we can train our entire model end-to-end.
However, to leverage more weakly-labeled data, we can also decompose Dense VOC into subtasks, and use each subtask to supervise the relevant part of our model using the available annotations as shown in Tab. 1. This approach also enables us to perform our final task in a zero-shot manner (i.e. without training on any full Dense VOC annotations).
Object detection. Using detection datasets for images, we can train the object proposal generator with $L_{object}$. We use COCO [41] as it is the most popular dataset for this task.
Dense captioning in images. Dense object captioning datasets of images allow us to train both the object proposal generator and the text decoder, by supervising $L_{object}$ and $L_{caption}$. Here, we use Visual Genome [34], the largest dataset for this task.
Global video captioning. Video captioning datasets help us to reduce the domain gap to our final task by also training on video. In particular, we use Spoken Moments in Time (SMiT) [45] which is the largest dataset for this task and contains narrations for short clips (roughly 3 seconds). As there are no object annotations, but only video-level captions, we construct an object proposal from the entire frame and caption that with our text decoder, applying $L_{caption}$. This approach is inspired by prior work on weakly-supervised object detection [86, 8, 4].
Tracking. Training the tracking module of our network (Sec. References) requires annotations that associate detections of an object identity throughout the video. We found that existing tracking datasets either have too limited vocabularies for general objects (MOT [19], KITTI [25], YouTube VIS [74]), or are too small (TAO [17] and UVO [63] label 600 and 5 000 videos respectively). As a result, following existing work [82, 89, 87], we instead augment image datasets into tracking ones by applying two different data augmentations to the same image, and then linearly interpolating the frames in between to form a pseudo-video. In particular, we augment COCO (referred to as Aug-COCO [87]). This enables us to apply $L_{assoc}$ and $L_{object}$ when training our model.
Dataset Annotation type Train size $L_{object}$ $L_{assoc}$ $L_{caption}$
COCO[41] Image detection 118K
VG[34] Image object captioning 70K
SMiT[45] Video captioning 480K
Aug-COCO[41] Video object tracking 118K
Table 1: Datasets for disjoint pretraining. We supervise different losses based on available annotations from each dataset.
@@ 3.5 Application to video object grounding
The task of video object grounding [83, 60] consists of two inputs: a video, $\mathbf{V}$, and a sentence query, $\bar{c}$. The output is a sequence of bounding boxes, $[b^{s},b^{s+1},\dots,b^{e}]$, corresponding to the sentence query, where $s$ and $e$ are the indices of the start and end frames respectively.
Our model, however, generates captions, $c$, at the output, rather than requiring it as an input. To apply our model to grounding, we follow an analogous approach to prior works that performed closed-set image classification with captioning models [1, 12]: we evaluate the likelihood (i.e., exponential negative cross-entropy loss) of the sentence query, $\bar{c}$, for each of the object trajectories produced by our model. In practice, we find that instead of just taking the object trajectory with the highest sentence-likelihood, we achieve higher accuracy by weighting the likelihood by the detection score, $s_{k}$, from our object proposal module. Thus, given bounding boxes, trajectory features and detection scores, $\{(b_{k}^{t},s_{k}^{t},\mathbf{g}_{k})\}_{t=1,k=1}^{T,K_{t}}$, we choose the bounding boxes with the highest weighted likelihood:
${k}^{*}$ $=\operatorname*{arg\,max}_{k}\left(s_{k}^{t}\cdot\text{exp}(-L_{caption}(\text{Decode}(\mathbf{f}_{k}^{t}),\bar{c}))\right),$
$b^{t}$ $=b_{{k}^{*}}^{t}.$ (1)
@@ 4 Experimental Evaluation
As we are proposing a new task, there is no dedicated dataset or evaluation metric for dense video object captioning. Fortunately, existing video grounding [83, 60] datasets have annotations for object trajectories and their captions, allowing us to repurpose them for Dense VOC, as defined next.
@@ 4.1 Datasets
VidSTG [83] was originally created for spatio-temporal sentence grounding, but can be used for Dense VOC: Each video annotates multiple textual queries and their corresponding spatio-temporal tubes. By aggregating these across all videos, we obtain the paired trajectory-caption annotations that we need for training and evaluating our model.
VidSTG has exhaustive trajectory (i.e. bounding box and tracking) annotations for all objects [53], but not all objects are used in grounding, and thus not all objects have captions. We account for this fact in both training and testing. Specifically, we do not compute $L_{caption}$ on objects without caption annotations, and also exclude them during evaluation (see Sec. 4.2). In particular, when a prediction is matched to a ground truth without caption annotations, we do not evaluate its captioning metrics, but still evaluate detection metrics. The dataset contains 5 436 training videos and 602 validation videos, with each video being at most 200 frames long. We use the declarative annotations from the dataset containing 19 000 captioned trajectories for training and testing respectively.
Video Localized Narratives (VLN) [60] augments existing datasets by narrating the “main actors” in a video. We therefore use these narrations as our target captions. We use the subset from the UVO dataset [63] as UVO has exhaustive detection and tracking annotations for all objects. Like VidSTG, the captions are not exhaustive for all objects, so we exclude objects without captions in both training and evaluating the captioning module. Each video has bounding box annotations for 3 sparsely sampled frames, and thus we train and evaluate on these frames. The dataset contains a total of 5 136 training and 2 451 validation videos, and 5 588 training and 3 071 validation captioned trajectories.
@@ 4.2 Evaluation metrics
[…]
which uses three popular image-captioning metrics [13], and $TP^{\prime}$ are the true-positive detection pairs that have caption annotations (as discussed in Sec. References). Note that for compatibility with HOTA, we follow DetA and AssA and compute CapA separately per-object on each frame. The final metric is then $\text{CHOTA}=\sqrt[3]{\text{DetA}\cdot\text{AssA}\cdot\text{CapA}}$, effectively adding a captioning term to the HOTA metric. Further details are in the supplement.
Frame mAP-METEOR ($\text{AP}_{M}$). We additionally reuse the evaluation metric for dense object captioning in images from Visual Genome [34]. Note that we avoid the captioning threshold for objects without captioning annotations, as detailed in the supplement. As the original metric was developed for images, it is computed separately on each frame.
@@ 4.3 Implementation details
We implement our model in JAX [9] using the scenic libinary [18]. We use GRiT [65] as our baseline. GRiT uses a ViTDet-Base [22, 39] backbone (initialized with CLIP pretrained parameters [48]), a CenterNet [88] object proposal network and RoI Head, and a randomly-initialized text decoder.
We first train our model for general Dense VOC on large-scale disjoint datasets (Sec. References). During disjoint pretraining, we sample batches from different datasets with an even ratio, (1: 1: 1: 1), thus avoiding additional hyperparameters. For video datasets, we sample 8 frames for a video and use a local batch size of 1. For image datasets, we use a local batch size of 8. We train our model on 32 TPUs, which means we have an effective batch size of 256 images or 32 videos.
We then evaluate the models on the two fully-annotated datasets (Sec. References) in both zero-shot and full-finetuning setups. For VidSTG, we sample 16 frames during training, and then run on all 200 frames during testing. For VLN, we use all 3 annotated frames in both training and evaluation. In both cases, we use an input size of $384\!\times\!384$. During inference, we threshold the outputs of our object proposal module with a score of $0.5$, and only track the remaining objects. We include exhaustive implementation details and hyperparameters in the supplement and the code.
4.4 Analysis of end-to-end tracking
@@
2 Per-frame cap. w. FairMOT [82] 51.2 63.4 57.2 37.0
3 Per-frame cap. w. ByteTrack [81] 52.3 64.2 60.2 37.1
4 Middle-frame cap. w. ByteTrack [81] 50.7 64.2 60.2 33.8
@@
Table 2: Comparison of our end-to-end model to per-task baselines on VidSTG validation. Our models are based on #2 of Tab. References. The image dense captioning models used in the baselines (rows #1-#4), are trained on the same datasets, and run off-the-shelf trackers as a post-processing step. Our end-to-end approach improves across all metrics, and additionally produces temporally consistent captions.
We first study the benefits of our end-to-end model, and the importance of captioning-after-tracking, in Tab. References. We do this by comparing to multiple, strong baseline models running in sequence. Concretely, we use the state-of-the-art image-dense object captioning model [65] followed by tracking as a post-processing step. We use trackers ranging from a simple IoU-based tracker [67] to more recent, sophisticated methods like FairMOT [82], and ByteTrack [81].
@@
As shown in Tab. References, per-frame captioners followed by offline trackers produce temporally inconsistent captions (#1-#3). Naively selecting the caption from the middle frame as the trajectory-level caption produces temporally consistent captions, but comes at the cost of captioning accuracy, as a single frame may not be representative of the entire event (#4). Both variants of our model (#5 and #6) improve tracking quality substantially, as shown by their large improvement on AssA, demonstrating the benefits of end-to-end training and incorporating temporal information. Our model improves on CapA too, showing that improved object trajectories provide better features for subsequent captioning. Finally, we note that as expected, the quality of the initial detections at each frame, measured by DetA, does not really change between the baselines and our method. This does, however, show that training our model jointly with multiple loss functions does not compromise performance on individual tasks.
Overall, our end-to-end model (#6) improves the CHOTA by 2.6 points over the best baseline (#3), due to our improvements in AssA and CapA. As hard aggregation performs slightly better, we use it in our following experiments.
@@ 4.5 Analysis of disjoint training
Zero-shot evaluation.
# COCO VG SMiT Aug-COCO VidSTG VLN
CHOTA DetA AssA CapA $\text{AP}_{M}$ CHOTA DetA AssA CapA $\text{AP}_{M}$
1 - 48.9 - - - - 27.8 - - -
2 - 17.8 - 7.8 17.1 - 12.1 - 7.4 9.9
3 - - - - - - - - - -
4 - 19.1 - 8.5 18.2 - 14.3 - 8.5 12.7
5 - 49.9 - 8.1 37.4 - 28.0 - 7.8 19.7
6 - 50.4 - 4.9 36.7 - 28.7 - 7.5 18.1
7 - 51.3 - 9.1 38.2 - 29.9 - 9.0 19.0
8 31.1 51.4 59.6 9.8 39.5 29.2 29.1 88.0 9.7 20.1
@@
Table 3: Zero-shot evaluation of our disjoint trained models with varying datasets. We show results on VidSTG [83] (center) and VLN [60] (right) without finetuning on them. For models without tracking supervision (#1–7), we cannot report their association accuracy (AssA). Our full model (#8) gains full Dense VOC ability from disjoint training, and shows good performance on all metrics. Noteably, CapA is improved given tracking supervision. Detailed captioning metrics are in the supplement.
@@
We first pretrain on multiple disjoint datasets (Sec. 3.4), and evaluate zero-shot on our target datasets, VidSTG and VLN, without training on them in Tab. References. Zero-shot evaluation is simple to perform for models that directly generate text compared to classification, thanks to their open vocabulary. There are, however, significant domain gaps in the object classes, captioning style, and video durations between our training and evaluation datasets.
@@
As mentioned in Tab. 1 and Fig. 3, each dataset supervises different parts of our model. For example, a model that is only trained on COCO (#1 in Tab. References), is only trained with $L_{object}$, meaning that it only produces object proposals which we can evaluate with the Detection Accuracy component of our CHOTA metric. Visual Genome (VG) can supervise both the object proposal and captioning heads of our model. However, there is a large domain gap between the captions in VG and our target datasets, since the captions in VG are for single images and focus on very different vocabularies. Furthermore, VG tends to annotate bounding boxes around object parts rather than entire objects. Consequently, our zero-shot DetA is low when training only on VG (#2). Note that we cannot evaluate a model trained only on SMiT, as it does not produce bounding boxes.
@@
We observe in Tab. References that the different datasets have complementary properties: Adding COCO improves detection accuracy (#2 to #5, #4 to #7), and adding SMiT improves the captioning accuracy (#2 to #4, #5 to #7) even though SMiT only captions at a video-level. Note that to mitigate the differences in the type of bounding boxes annotated by VG, we ignore $L_{object}$ on it and only compute $L_{caption}$, when using it in conjunction with COCO. Finally, training with Aug-COCO allows us to also supervise $L_{assoc}$ and thus the tracking module of our model. A model trained on all the datasets (#8) can therefore perform the full Dense VOC task, and shows good performance on all individual metrics compared to models trained on fewer datasets. Notably, we observe our final model with tracking improves captioning ability (CapA) without adding captioning training data. Again, similar to Tab. References, the improvements are likely to result from our ability to leverage temporal information from tracking.
# COCO VG SMiT Aug-COCO VidSTG VLN
CHOTA DetA AssA CapA $\text{AP}_{M}$ CHOTA DetA AssA CapA $\text{AP}_{M}$
0 47.8 54.6 57.8 34.5 54.1 29.7 35.3 85.4 8.7 35.1
1 52.3 64.9 63.0 34.9 69.1 31.8 43.9 88.7 8.2 36.3
2 54.9 64.2 65.9 39.1 68.7 40.6 45.1 88.4 16.7 45.9
3 45.4 51.9 56.9 31.6 54.8 37.4 41.2 87.7 14.5 38.0
4 55.2 64.0 67.1 39.2 68.9 41.0 44.2 88.4 17.8 47.2
5 55.6 65.7 68.9 38.4 70.8 40.9 44.1 88.8 17.4 46.1
6 54.4 64.9 63.9 38.8 69.4 35.6 43.7 88.5 11.6 41.3
7 56.5 65.8 68.2 40.1 71.2 41.1 44.2 88.9 17.7 48.2
8 56.9 65.8 70.4 39.7 71.5 41.3 44.3 89.5 17.7 48.2
Table 4: Finetuning evaluation of our disjoint trained models with varying pretraining datasets. We show results on VidSTG [83] (center) and VLN [60] (right). Each row is a model pretrained on the specified datasets and then finetuned on the downstream datasets. #0 is finetuned from a CLIP [48] checkpoint. Our disjoint-pretrained models (#${\color[rgb]{0.4453125,0.625,0.8125}\#4}\!-\!{\color[rgb]{0.4453125,0.625,0.8125}\#8}$) perform better than single-dataset pretrained ones (#${\color[rgb]{0.4453125,0.625,0.8125}\#0}\!-\!{\color[rgb]{0.4453125,0.625,0.8125}\#3}$), and yield the best CHOTA and $\text{AP}_{M}$ on both VidSTG and VLN. The best entries are bolded, and the second-best underlined. Detailed caption metrics are in the supplement.

### v3

@@ 3.3 Trajectory captioning
Soft aggregation. A straightforward way to leverage object features over time is to compute a weighted sum to combine them into a single, global trajectory feature. We observe that the association matrix, $\mathbf{A}$ (Sec. 3.2), already serves as a summation weight. Specifically, we set $\mathbf{G}=\frac{\mathbf{A}}{||\mathbf{A}||}\cdot\mathbf{F}$, where $\cdot$ denotes matrix multiplication, and |||||$\cdot$||| | ⋅ | | normalizes $\mathbf{A}$ by rows. Each row of $\mathbf{G}$, $\mathbf{g}_{k}\in\mathbb{R}^{r^{2}\times d}$, therfore denotes an aggregated feature over its trajectory for object $k$.
Hard aggregation. An alternative to weighted temporal averaging is to concatenate and construct new trajectory features. Let $\mathbf{f}_{\tau}=\{\mathbf{f}_{k^{\prime}}\}_{\delta_{k^{\prime}}=\tau}$ be the set of all object features with identity $\tau$. We note $\mathbf{f}_{\tau}$ can be as long as the entire video, and thus it may be expensive to directly use $\mathbf{f}_{\tau}$. Therefore, we uniformly sample a subset of object features from the trajectory, denoted as $\mathbf{g}_{\tau}=\text{UniformSample}(\mathbf{f}_{\delta},m)$, where $\mathbf{g}_{\tau}\in\mathbb{R}^{mr^{2}\times d}$, inspired by Wang et al. (2022). $m$ is the number of sampled frames, and we set $m=6$ following ablations in Appendix C.2.
@@ 3.4 Pretraining with disjoint subtasks
As shown in Fig. 3, our model is trained with the loss function, $L=L_{object}+L_{assoc}+L_{caption}$. When we have full Dense VOC annotations, which supervise each component of our model we can train our entire model end-to-end. However, to leverage more weakly-labeled data, we can also decompose Dense VOC into subtasks, and use each subtask to supervise the relevant part of our model using the available annotations as shown in Tab. 1. This approach also enables us to perform our final task in a zero-shot manner (i.e. without training on any full Dense VOC annotations).
Dataset Annotation type Train set size ($10^{3}$) $L_{object}$ $L_{assoc}$ $L_{caption}$
COCO (Lin et al., 2014) Image detection 118
VG (Krishna et al., 2017b) Image object captioning 70
SMiT (Monfort et al., 2021) Video captioning 480
Aug-COCO (Lin et al., 2014) Video object tracking 118
Table 1: Datasets for pretraining. We supervise different losses based on available annotations.
Object detection. Using detection datasets for images, we can train the object proposal generator with $L_{object}$. We use COCO (Lin et al., 2014) as it is the most popular dataset for this task.
Dense captioning in images. Dense object captioning datasets of images allow us to train both the object proposal generator and the text decoder, by supervising $L_{object}$ and $L_{caption}$. Here, we use Visual Genome (Krishna et al., 2017b), the largest dataset for this task.
Global video captioning. Video captioning datasets help us to reduce the domain gap to our final task by also training on video. In particular, we use Spoken Moments in Time (SMiT) (Monfort et al., 2021) which is the largest dataset for this task and contains narrations for short clips (roughly 3 seconds). As there are no object annotations, but only video-level captions, we construct an object proposal from the entire frame and caption that with our text decoder, applying $L_{caption}$. This approach is inspired by prior work on weakly-supervised object detection (Zhou et al., 2022a; Bilen & Vedaldi, 2016; Arnab et al., 2020).
Tracking. Training the tracking module of our network (Sec. 3.2) requires annotations that associate detections of an object identity throughout the video. We found that existing tracking datasets either have too limited vocabularies for general objects (MOT (Dendorfer et al., 2021), KITTI (Geiger et al., 2012), YouTube VIS (Yang et al., 2019)), or are too small (TAO (Dave et al., 2020) and UVO (Wang et al., 2021b) label 600 and 5 000 videos respectively), and thus giving unsatisfactory results in our setting (Appendix C.3). As a result, following existing work (Zhang et al., 2021b; Zhou et al., 2020), we instead augment image datasets into tracking ones by applying two different data augmentations to the same image, and then linearly interpolating the frames in between to form a pseudo-video. In particular, we augment COCO (referred to as Aug-COCO (Zhou et al., 2020)). This enables us to apply $L_{assoc}$ and $L_{object}$ when training our model.
@@ 3.5 Application to video object grounding
The task of video object grounding (Zhang et al., 2020; Voigtlaender et al., 2023) consists of two inputs: a video, $\mathbf{V}$, and a sentence query, $\bar{c}$. The output is a sequence of bounding boxes, $[b^{s},b^{s+1},\dots,b^{e}]$, corresponding to the sentence query, where $s$ and $e$ are the indices of the start and end frames respectively.
Our model, however, generates captions, $c$, at the output, rather than requiring it as an input. To apply our model to grounding, we follow an analogous approach to prior works that performed closed-set image classification with captioning models (Alayrac et al., 2022; Chen et al., 2023b): we evaluate the likelihood (i.e., exponential negative cross-entropy loss) of the sentence query, $\bar{c}$, for each of the object trajectories produced by our model. In practice, we find that instead of just taking the object trajectory with the highest sentence-likelihood, we achieve higher accuracy by weighting the likelihood by the detection score, $s_{k}$, from our object proposal module. Thus, given bounding boxes, trajectory features and detection scores, $\{(b_{k}^{t},s_{k}^{t},\mathbf{g}_{k})\}_{t=1,k=1}^{T,K_{t}}$, we choose the bounding boxes with the highest weighted likelihood:
${k}^{*}=\operatorname*{arg\,max}_{k}\left(s_{k}^{t}\cdot\text{exp}(-L_{caption}(\text{Decode}(\mathbf{f}_{k}^{t}),\bar{c}))\right),\qquad b^{t}=b_{{k}^{*}}^{t}.$ (1)
@@ 4 Experimental Evaluation
As we are proposing a new task, there is no dedicated dataset or evaluation metric for dense video object captioning for all objects. Fortunately, existing video grounding datasets (Zhang et al., 2020; Voigtlaender et al., 2023) have annotations for object trajectories and their captions, allowing us to repurpose them for Dense VOC, as defined next. We also report results on concurrent person-focused video object tracking and captioning benchmark, BenSMOT(Li et al., 2024b).
@@ 4.1 Datasets
VidSTG (Zhang et al., 2020) was originally created for spatio-temporal sentence grounding, but can be used for Dense VOC: Each video annotates multiple textual queries and their corresponding spatio-temporal tubes. By aggregating these across all videos, we obtain the paired trajectory-caption annotations that we need for training and evaluating our model.
VidSTG has exhaustive trajectory (i.e. bounding box and tracking) annotations for all objects (Shang et al., 2019), but not all objects are used in grounding, and thus not all objects have captions. We account for this fact in both training and testing. Specifically, we do not compute $L_{caption}$ on objects without caption annotations, and also exclude them during evaluation (see Sec. 4.2). In particular, when a prediction is matched to a ground truth without caption annotations, we do not evaluate its captioning metrics, but still evaluate detection metrics. The dataset contains 5,436 training videos and 602 validation videos, with each video being at most 200 frames long. We use the declarative annotations from the dataset containing 19,000 captioned trajectories for training.
Video Localized Narratives (VLN) (Voigtlaender et al., 2023) augments existing datasets by narrating the “actors” in a video. We therefore use these narrations as our target captions. We use the subset from the UVO dataset (Wang et al., 2021b) as UVO has exhaustive detection and tracking annotations for all objects. Like VidSTG, the captions are not exhaustive for all objects, so we exclude objects without captions in both training and evaluating the captioning module. Each video has bounding box annotations for 3 sparsely sampled frames, and thus we train and evaluate on these frames. The dataset contains a total of 5,136 training and 2,451 validation videos.
BenSMOT (Li et al., 2024b) contains person bounding boxes trajectories and their manually-annotated captions for 3292 YouTube videos. The dataset has in average 2.2 trajectories per video.
@@ 4.2 Evaluation metrics
$\text{CapA}=\frac{1}{3|TP^{\prime}|}\sum_{(x,y)\in TP^{\prime}}({\text{METEOR}(x,y)+\text{CIDEr}(x,y)+\text{SPICE}(x,y)}),$ (2)
which uses three popular image-captioning metrics (Chen et al., 2015), and $TP^{\prime}$ are the true-positive detection pairs that have caption annotations (as discussed in Sec. 4.1). Note that for compatibility with HOTA, we follow DetA and AssA and compute CapA separately per-object on each frame. The final metric is then $\text{CHOTA}=\sqrt[3]{\text{DetA}\cdot\text{AssA}\cdot\text{CapA}}$, effectively adding a captioning term to the HOTA metric. We include further details and code in Appendix B, along with results using the image dense object captioning metrics, mAP-METEOR (Appendix C.1).
@@ 4.3 Implementation details
Our implementation is based on the public release of GRiT (Wu et al., 2022a). GRiT uses a ViTDet-Base (Li et al., 2022b) backbone (initialized with CLIP (Radford et al., 2021)), a CenterNet (Zhou et al., 2019) object proposal network and RoI Head, and a randomly-initialized text decoder.
We first train our model for general Dense VOC on large-scale disjoint datasets (Sec. 3.4). During disjoint pretraining, we sample batches from different datasets with an even ratio, (1: 1: 1: 1), thus avoiding additional hyperparameters. For video datasets, we sample 8 frames for a video and use a local batch size of 1. For image datasets, we use a local batch size of 8. We train our model on 32 GPUs, which means we have an effective batch size of 256 images or 32 videos.
@@
2 Per-frame cap. w. FairMOT Zhang et al. (2021b) 51.2 63.4 57.2 37.0
3 Per-frame cap. w. ByteTrack Zhang et al. (2022b) 52.3 64.2 60.2 37.1
4 Middle-frame cap. w. ByteTrack Zhang et al. (2022b) 50.7 64.2 60.2 33.8
@@
Table 2: Comparison of our end-to-end model to per-task baselines on VidSTG validation. Our models are based on #2 of Tab. 3 right. The image dense captioning models used in the baselines (rows #1-#4) are trained on the same datasets, and run off-the-shelf trackers as post-processing. Our end-to-end approach improves across all metrics, and produces temporally consistent captions.
We then evaluate the models on the two fully-annotated datasets (Sec. 4.1) in both zero-shot and full-finetuning setups. For VidSTG, we sample 16 frames during training, and then run on all 200 frames during testing. For VLN, we use all 3 annotated frames in both training and evaluation. In both cases, we use an input size of $384\!\times\!384$. During inference, we threshold the outputs of our object proposal module with a score of $0.5$, and only track the remaining objects. We include exhaustive implementation details and hyperparameters in Appendix B.2 with the full code.
4.4 Analysis of end-to-end tracking
We first study the benefits of our end-to-end model in Tab. 2. We do this by comparing to multiple, strong baseline models running in sequence. Concretely, we use the state-of-the-art image-dense object captioning model  Wu et al. (2022a) followed by tracking as a post-processing step. We use trackers ranging from a simple IoU-based tracker Wu et al. (2019) to more recent, sophisticated methods like FairMOT (Zhang et al., 2021b), and ByteTrack (Zhang et al., 2022b).
@@ 4.4 Analysis of end-to-end tracking
As shown in Tab. 2, per-frame captioners followed by offline trackers produce temporally inconsistent captions (#1-#3). Naively selecting the caption from the middle frame as the trajectory-level caption produces temporally consistent captions, but comes at the cost of captioning accuracy, as a single frame may not be representative of the entire event (#4). Both variants of our model (#5 and #6) improve tracking quality substantially, as shown by their large improvement on AssA, demonstrating the benefits of end-to-end training and incorporating temporal information. Our model improves on CapA too, showing that improved object trajectories provide better features for subsequent captioning. Finally, we note that the quality of the initial detections at each frame, measured by DetA, does not really change between the baselines and our method. This does, however, show that training our model jointly with multiple loss functions does not compromise performance on individual tasks.
Overall, our end-to-end model (#6) improves the CHOTA by 2.6 points over the best baseline (#3). As hard aggregation performs slightly better, we use it in our following experiments.
@@ 4.5 Analysis of disjoint training
# COCO VG SMiT Aug VidSTG (zero-shot) VLN (zero-shot) VidSTG (finetuned) VLN (finetuned)
COCO CHOTA DetA AssA CapA CHOTA DetA AssA CapA CHOTA DetA AssA CapA CHOTA DetA AssA CapA
0 - - - - - - - - 47.8 54.6 57.8 34.5 29.7 35.3 85.4 8.7
1 - 48.9 - - - 27.8 - - 52.3 64.9 63.0 34.9 31.8 43.9 88.7 8.2
2 - 17.8 - 7.8 - 12.1 - 7.4 54.9 64.2 65.9 39.1 40.6 45.1 88.4 16.7
3 - - - - - - - - 45.4 51.9 56.9 31.6 37.4 41.2 87.7 14.5
4 - 19.1 - 8.5 - 14.3 - 8.5 55.2 64.0 67.1 39.2 41.0 44.2 88.4 17.8
5 - 49.9 - 8.1 - 28.0 - 7.8 55.6 65.7 68.9 38.4 40.9 44.1 88.8 17.4
6 - 50.4 - 4.9 - 28.7 - 7.5 54.4 64.9 63.9 38.8 35.6 43.7 88.5 11.6
7 - 51.3 - 9.1 - 29.9 - 9.0 56.5 65.8 68.2 40.1 41.1 44.2 88.9 17.7
8 31.1 51.4 59.6 9.8 29.2 29.1 88.0 9.7 56.9 65.8 70.4 39.7 41.3 44.3 89.5 17.7
@@
Table 3: Zero-shot (left) and finetuning (right) evaluation of our disjoint trained models with varying datasets. We show results on VidSTG (Zhang et al., 2020) and VLN (Voigtlaender et al., 2023). Each row is a model pretrained on the specified datasets for zero-shot evaluation and then finetuned on the downstream datasets. #0 is finetuned from a CLIP checkpoint. For models without tracking supervision (#1–7), we cannot report their zero-shot association accuracy (AssA). Our full model (#8) gains full Dense VOC ability from disjoint training, and shows good performance on all metrics with or without finetuning, on both datasets. Detailed captioning metrics are in Appendix C.4.
@@
Zero-shot evaluation. We first pretrain on multiple disjoint datasets (Sec. 3.4), and evaluate zero-shot on our target datasets, VidSTG and VLN, without training on them in Tab. 3 left. Zero-shot evaluation is simple to perform for captioning models compared to classification, thanks to their open vocabulary.
@@
As mentioned in Tab. 1 and Fig. 3, each dataset supervises different parts of our model. For example, a model that is only trained on COCO (#1 in Tab. 3), is only trained with $L_{object}$, meaning that it only produces object proposals which we can evaluate with the Detection Accuracy component of our CHOTA metric. Visual Genome (VG) can supervise both the object proposal and captioning heads of our model. However, there is a large domain gap between the captions in VG and our target datasets, since the captions in VG are for single images and focus on very different vocabularies. Furthermore, VG tends to annotate bounding boxes around object parts rather than entire objects. Consequently, our zero-shot DetA is low when training only on VG (#2). To mitigate the differences in the type of bounding boxes annotated by VG, we ignore $L_{object}$ on it when using it in conjunction with COCO. Note that we cannot evaluate a model trained only on SMiT, as it does not produce bounding boxes.
@@
We observe in Tab. 3 (left) that the different datasets have complementary properties: Adding COCO improves detection accuracy (#2 to #5, #4 to #7), and adding SMiT improves the captioning accuracy (#2 to #4, #5 to #7) even though SMiT only captions at a video-level. Finally, training with Aug-COCO allows us to also supervise $L_{assoc}$ and thus the tracking module of our model. A model trained on all the datasets (#8) can therefore perform the full Dense VOC task, and shows good performance on all individual metrics compared to models trained on fewer datasets. Notably, we observe our final model with tracking improves captioning ability (CapA) without adding captioning training data. Similar to Tab. 2, the improvements are likely from our ability to leverage temporal information.